# Simon Yu1  Gang Li2  Weiyan Shi†1  Peng Qi†2

### v1

@@ Abstract
Large language models (LLMs) are moving beyond static uses and are now powering agents that learn continually during their interaction with external environments. For example, agents can learn reusable skills while navigating web pages or toggling new tools. However, existing methods for skill learning often create skills that are over-specialized to a single website and fail to generalize. We introduce PolySkill, a new framework that enables agents to learn generalizable and compositional skills. The core idea, inspired by polymorphism in software engineering, is to decouple a skill’s abstract goal (what it accomplishes) and its concrete implementation (how it is executed). Experiments show that our method (1) improves skill reuse by 1.7x on seen websites and (2) boosts success rates by up to 9.4% on Mind2Web and 13.9% on unseen websites, while reducing steps by over 20%. (3) In self-exploration settings without specified tasks, our framework improves the quality of proposed tasks and enables agents to learn generalizable skills that work across different sites. By enabling the agent to identify and refine its own goals, the PolySkill enhance the agent a better curriculum, leading to the acquisition of more generalizable skills compared to baseline methods. This work provides a practical path toward building agents capable of continual learning in adaptive environments. Our findings show that separating a skill’s goal from its execution is a crucial step toward developing autonomous agents that can learn and generalize across the open web continuously.
@@ 1 Introduction
LLMs have enabled significant progress in agent development (Yao et al., 2023b; Yang et al., 2024; Agashe et al., 2025). Web agents represent a key class of LLM-based agents, where the agents navigates complex Graphical User Interfaces (GUIs) to achieve user-defined goals (Deng et al., 2023; Zhou et al., 2024a; Cheng et al., 2024; Wu et al., 2024b; Chen et al., 2025b; Gou et al., 2025; Xue et al., 2025). However, a primary challenge here is how to develop generalizable agents that can operate robustly across different tasks within a website (in-domain) and across different websites (cross-domain). To be generalizable, these agents must learn from their experiences, allowing them to adapt when faced with new tasks or unseen websites (Silver & Sutton, 2025; Shen et al., 2025).
One promising direction is skill induction: learning reusable skills from past experiences. This approach was first explored in an open-ended environment by Voyager (Wang et al., 2023). Agent Workflow Memory (Wang et al., 2024) pioneered skill induction for web agents, which used natural language skills to prove the concept’s viability. Consequently, Agent Skill Induction (Wang et al., 2025) and SkillWeaver (Zheng et al., 2025) made these skills more robust by structuring them as code. However, these methods primarily focus on in-domain settings. By optimizing for performance on familiar websites, they generate over-specialized skills that fail to generalize, leaving the critical challenge of cross-domain generalization under-explored. This challenge highlights a fundamental tension between a skill’s specificity and its generalizability. This leads us to two key research questions. First, how can we induce skills that are transferable across diverse websites? Second, beyond task success, how can we quantitatively measure skill transfer and reuse? Answering these questions is a crucial step toward the broader vision of agent autonomy, where an agent could discover such generalizable skills on its own through open-ended exploration (Liu et al., 2025a).
@@ 2 Related Work
Memory and Skill Acquisition for Agentic Learning An agent’s ability to generalize is fundamentally tied to how it represents skills. Current approaches often store skills in concrete formats, such as natural language descriptions in prompt libraries (Wang et al., 2024; Zhu et al., 2025) or brittle action traces from successful task executions (Wang et al., 2025).The idea of agent skills is also recently used by Anthropic, which allows users to provide a curated set of tools for the agent to invoke (Anthropic, 2025), but this still requires manual effort in defining such skills. While more robust, programmatic representations, where skills are stored as executable code, also face limitations. These learned programs are typically concrete implementations tailored to a specific context (e.g., one website’s UI), hindering their reuse across varied environments that serve a similar function (Wang et al., 2023; Zheng et al., 2025). These methods lack a mechanism for abstracting the semantic intent of a skill away from its specific implementation, which is crucial for flexible adaptation.
@@
We address this gap with a novel, hybrid skill representation inspired by object-oriented design. While broader work has focused on the architecture of agent memory, such as using episodic streams (Park et al., 2023) or managed external stores (Chhikara et al., 2025; Fang et al., 2025), our contribution lies in the representation of the skill itself. We define a skill with an abstract interface that captures its semantic purpose, which can be linked to multiple, interchangeable concrete implementations. Drawing on established findings that abstraction and structure enhance agent reasoning and robustness (Wu et al., 2024a; Yao et al., 2023a), this polymorphic structure provides the formal benefits of programmatic skills while explicitly incorporating the flexibility required for generalization across different domains.
@@
Continual Learning A long-standing goal in AI is continual learning (van de Ven et al., 2025; Liu et al., 2025a), where an agent continually learn during testing time and its interaction with the environments. This is specifically true for web agents, where interaction is more effective than just thinking longer (Shen et al., 2025; Liu et al., 2025b). A key insight from AI research is that this requires compositional generalization, the ability to learn primitive concepts and recombine them to solve novel problems (Jiang et al., 2025). Existing web agents that learn from experience, such as those using exploration enhanced by curriculum learning (Zheng et al., 2025) or reinforcement learning (Zhou et al., 2024b), perform a simple form of continual learning by adding new skills to a library. However, their skills lack a compositional structure; a learned code snippet or action trace cannot be easily modified or combined with others in a principled way, leading to low utilization rates on new websites. Our work, first showing the effectiveness of the principles for modularity from software engineering, offers a powerful solution. Polymorphism (Milner, 1978) is a time-tested concept designed explicitly to manage variation between implementations while maintaining a stable interface. By applying this principle to agent skills, our approach provides a structured paradigm for learning capabilities that are modular and interchangeable. See Appendix A for detailed related work.
@@ 3.1 Preliminary
Problem Formulation. We model the web agent’s interaction environment as a Partially Observable Markov Decision Process (POMDP), defined by the tuple $\langle\mathcal{S},\mathcal{A}_{p},\mathcal{T},\Omega,\mathcal{O}\rangle$. Here, $\mathcal{S}$ is the latent state space representing the full underlying state of the web application. $\mathcal{A}_{p}$ is the set of primitive actions the agent can execute on a webpage (e.g., click(element), type(text)), as defined in Section C.2. The function $\mathcal{T}:\mathcal{S}\times\mathcal{A}_{p}\to\Delta(\mathcal{S})$ is the stochastic state transition function. Since the agent cannot perceive the entire state $\mathcal{S}$, it receives an observation $o_{t}\in\Omega$ (e.g., the A11y tree and viewport screenshot) at each timestep $t$ through the observation function $\mathcal{O}:\mathcal{S}\to\Delta(\Omega)$.
@@ 3.2 The PolySkill framework
We identified the limitations of current skill induction methods. We tested two state-of-the-art approaches, ASI (Wang et al., 2025) and SkillWeaver (Zheng et al., 2025), on how well their learned skills transfer to unseen websites. Their example of skills can be seen in . As illustrated in Figure 2, our analysis revealed two key problems. First, the learning process can be unstable, producing over-specialized skills. For instance, SkillWeaver’s performance with Claude-3.7-Sonnet degrades over time because its self-proposed tasks become increasingly complex and specific. This causes the resulting skills to be too intricate and poorly suited for generalization. Second, these skills show poor generalization when applied to new websites. This is reflected in extremely low Skill Reusabilitys on unseen websites: less than 9% for ASI and less than 3% for SkillWeaver.
We build on prior work that demonstrates the robustness of representing skills as code (Wang et al., 2025; Liu et al., 2025a). To specifically address the brittleness and over-specialization, we introduce a solution inspired from software engineering: Polymorphism (Milner, 1978). This allows our framework to separate a skill’s abstract goal from its concrete, site-specific implementation.
@@ 3.3 Skill Induction Process
Base of Skill Induction Pipeline Our skill induction process is built upon the robust verification pipeline established by ASI (Wang et al., 2025). In their framework, skill creation begins after a task is successfully completed using a sequence of primitive actions. An LLM-based induction module analyzes this successful trajectory to propose one or more programmatic skills that encapsulate reusable parts of the workflow (Pan et al., 2024). Before the skills are added into the library, a verification phase is done where the agent attempts to solve the same task again, this time by executing the newly generated skill. Only if this new execution is deemed as successful is the skill considered validated and added to the agent’s library for future use.
@@
Innovation via Polymorphic Skill Induction Where our method, PolySkill, improves is by integrating this process with a polymorphic skill structure. A critical preliminary step in our framework is that if the agent is operating on its first shopping website, it must first induce the high-level abstract class, AbstractShoppingSite, which defines a common skill interface for that category, including functions like search(item) and add_to_cart(item). Subsequently, when the agent generates a successful trajectory on a specific site like amazon.com, it begins to induce a concrete skill. We provide the AbstractShoppingSite class as context. This reframes the induction prompt: rather than asking the LLM to simply search for reusable skills, we instruct it to implement the specific search method for an AmazonWebsite class, ensuring it conforms to the abstract interface. This encourages the agent to learn skills that are not just locally effective but are structurally consistent implementations of a shared, domain-wide concept.
@@
Skill Learning on Unseen Websites This polymorphic structure makes learning on new websites within a known category significantly more efficient. Imagine the agent has already formed the AbstractShoppingSite class and now visits walmart.com for the first time. It immediately recognizes Walmart as a shopping site and retrieves the abstract blueprint. This blueprint gives the agent a clear set of goals for exploration. Instead of randomly trying actions, it knows it needs to figure out how to perform tasks like search_product and add_to_cart on this new site. Once the agent successfully searches for an item, it follows the standard induction process to create a new WalmartWebsite class, filling in the search method with the specific actions that worked. This guided approach accelerates learning by focusing the agent’s efforts on mastering the essential skills defined in the abstract interface.
@@ 3.4 Evaluation Setup
We evaluate our induction process over baseline, in two different settings: (1) Task-Defined Benchmarks In standard benchmark settings, we apply this process within controlled environments, including Mind2Web (Deng et al., 2023) and WebArena (Zhou et al., 2024a). Here, the agent is presented with a predefined curriculum of tasks. Each successful trajectory provides the validated sequence of actions needed to implement a concrete skill method. This controlled approach allows us to rigorously measure how well the polymorphic representation facilitates generalization and transfer to unseen tasks, websites, or domains. (2) Task-Free Continual Learning: To assess the ultimate goal of agent autonomy, we also apply our framework in a task-free setting, similar to settings as Voyager (Wang et al., 2023) and SkillWeaver (Zheng et al., 2025). In this scenario, the agent explores websites on its own, proposes its own goals, and induces skills from its successful attempts. Crucially, we showed that our polymorphic hierarchy enables structured exploration (Murty et al., 2025; Gandhi & Neubig, 2025). The already-learned abstract domain classes act as a schema, providing a strong prior for what skills are worth discovering (Liang et al., 2025).
@@ 3.5 Evaluation Metrics
We assess performance using five key metrics. Two are standard benchmarks adopted from prior work (Wang et al., 2024; 2025), while we introduce three new metrics to measure skills’ utility. (1) Task Success Rate (SR) is the percentage of held-out tasks the agent completes successfully. This is the primary measure of overall performance. (2) Number of Steps is the average number of actions the agent takes to complete a task, where each primitive action and each call to a skill is counted as a single step. Fewer steps indicate higher efficiency. (3) Skill Reusability measures the number skill reused in new tasks. A high rate indicates the agent learns relevant, broadly applicable skills rather than overly niche ones. (4) Task Coverage measures the tasks that used at least one skill. This metric indicates whether the skills can be adaptive in actual test case scenarios. (5) Skill Compositionality measures how frequently the system reuses existing skills as building blocks for more complex tasks. A high score indicates an efficient and scalable learning process, as the system leverages its acquired knowledge rather than learning every new skill from scratch. Formal definitions for all metrics are provided in Appendix C.1.
@@ 4.1 Standard Benchmark Evaluation
Results. For Mind2Web, as shown in Figure 3 (Results for WebArena is delayed in Figure 7). Our PolySkill outperforms ASI across both static and online settings. On GPT-4.1, PolySkill improves Cross-task accuracy from 52–53% (ASI) to 55–56%, and its online variant jumps to about 63%, compared to only 59% for ASI (+Update). In the hardest Cross-Domain setting, PolySkill (+Update) reaches 63–64% versus 62% for ASI (+Update). On Qwen3-Coder, PolySkill (+Update) improves Cross-task from 41.5% (ASI) to 47.5% and Cross-Domain from 35.2% to 39.9%. These consistent gains, particularly the significant jumps in Cross-Website, underscore PolySkill’s stronger generalization and the clear benefit of continual online updating over ASI.
Figure 3: Performance comparison of PolySkill with baseline methods on the Mind2Web benchmark across four large language models. The y-axis shows task success rate (%). The three evaluation settings on the x-axis, Cross-task, Cross-Website, and Cross-Domain, represent increasing levels of generalization difficulty. Our method, PolySkill, consistently outperforms the ASI baseline, while the online continual update version, PolySkill (+Online), achieves state-of-the-art performance across all models and settings. The performance gains are most significant in the challenging Cross-Domain scenario, highlighting our method’s superior generalization. Error bars represent the standard error over three runs.

### v2

@@ Abstract
Large language models (LLMs) are moving beyond static uses and are now powering agents that learn during their interaction with external environments. For example, agents can learn reusable skills while navigating web pages or toggling new tools. However, existing methods for skill learning often create skills that are over-specialized to a single website and fail to generalize. We introduce PolySkill, a new framework that enables agents to learn generalizable and compositional skills. The core idea, inspired by polymorphism in software engineering, is to decouple a skill’s abstract goal (what it accomplishes) and its concrete implementation (how it is executed). Experiments show that our method (1) improves skill reuse by 1.7x on seen websites and (2) boosts success rates by up to 9.4% on Mind2Web and 13.9% on unseen websites, while reducing steps by over 20%. (3) In self-exploration settings without specified tasks, our framework improves the quality of proposed tasks and enables agents to learn generalizable skills that work across different sites. By enabling the agent to identify and refine its own goals, the PolySkill enhance the agent a better curriculum, leading to the acquisition of more generalizable skills compared to baseline methods. This work provides a practical path toward building agents capable of continual learning in adaptive environments. Our findings show that separating a skill’s goal from its execution is a crucial step toward developing autonomous agents that can learn and generalize across the open web continuously. Our code can be found in https://github.com/simonucl/PolySkill.
@@ 1 Introduction
LLMs have enabled significant progress in agent development (Yao et al., 2023b; Yang et al., 2024; Agashe et al., 2025). Web agents represent a key class of LLM-based agents, where the agents navigates complex Graphical User Interfaces (GUIs) to achieve user-defined goals (Deng et al., 2023; Zhou et al., 2024a; Cheng et al., 2024; Wu et al., 2024b; Chen et al., 2025b; Gou et al., 2025; Xue et al., 2025). However, a primary challenge lies in developing generalizable agents that can operate robustly across different tasks within a single website, as well as transfer skills to distinct websites within the same functional domain (e.g., generalizing across different airline booking interfaces). To be generalizable, these agents must learn from their experiences, allowing them to adapt when faced with new tasks or unseen websites (Silver and Sutton, 2025; Shen et al., 2025).
One promising direction is skill induction: learning reusable skills from past experiences. This approach was first explored in an open-ended environment by Voyager (Wang et al., 2023). Agent Workflow Memory (Wang et al., 2024) pioneered skill induction for web agents, which used natural language skills to prove the concept’s viability. Consequently, Agent Skill Induction (Wang et al., 2025) and SkillWeaver (Zheng et al., 2025) made these skills more robust by structuring them as code. However, these methods primarily focus on same website, cross-task settings. By optimizing for performance on familiar websites, they generate over-specialized skills that fail to generalize, leaving the critical challenge of cross-website generalization under-explored. This challenge highlights a fundamental tension between a skill’s specificity and its generalizability. This leads us to two key research questions. First, how can we induce skills that are transferable across diverse websites? Second, beyond task success, how can we quantitatively measure skill transfer and reuse? Answering these questions is a crucial step toward the broader vision of agent autonomy, where an agent could discover such generalizable skills on its own through open-ended exploration (Liu et al., 2025a).
@@ 2 Related Work
Memory and Skill Acquisition for Agentic Learning An agent’s ability to generalize is fundamentally tied to how it represents skills. Current approaches often store skills in concrete formats, such as natural language descriptions in prompt libraries (Wang et al., 2024; Zhu et al., 2025) or brittle action traces from successful task executions (Wang et al., 2025). While more robust, programmatic representations, where skills are stored as executable code, also face limitations. These learned programs are typically concrete implementations tailored to a specific context (e.g., one website’s UI), hindering their reuse across varied environments that serve a similar function (Wang et al., 2023; Zheng et al., 2025). These methods lack a mechanism for abstracting the semantic intent of a skill away from its specific implementation, which is crucial for flexible adaptation.
@@
We address this gap with a novel, hybrid skill representation inspired by object-oriented design. While broader work has focused on the architecture of agent memory, such as using episodic streams (Park et al., 2023) or managed external stores (Chhikara et al., 2025; Fang et al., 2025), our contribution lies in the representation of the skill itself. We define a skill with an abstract interface that captures its semantic purpose, which can be linked to multiple, interchangeable concrete implementations. Drawing on established findings that abstraction and structure enhance agent reasoning and robustness (Wu et al., 2024a; Yao et al., 2023a), this polymorphic structure provides the formal benefits of programmatic skills while explicitly incorporating the flexibility required for generalization across different web pages.
@@
Continual Learning A long-standing goal in AI is continual learning (van de Ven et al., 2025; Liu et al., 2025a), where an agent continually learn during testing time and its interaction with the environments. This is specifically true for web agents, where interaction is more effective than just thinking longer (Shen et al., 2025; Liu et al., 2025b). A key insight from AI research is that this requires compositional generalization, the ability to learn primitive concepts and recombine them to solve novel problems (Jiang et al., 2025). Existing web agents that learn from experience, such as those using exploration enhanced by curriculum learning (Zheng et al., 2025) or reinforcement learning (Zhou et al., 2024b), perform a simple form of continual learning by adding new skills to a library. However, their skills lack a compositional structure; a learned code snippet or action trace cannot be easily modified or combined with others in a principled way, leading to low utilization rates on new websites. Our work, first showing the effectiveness of the principles for modularity from software engineering, offers a powerful solution. Polymorphism (Milner, 1978) is a time-tested concept designed explicitly to manage variation between implementations while maintaining a stable interface. By applying this principle to agent skills, our approach provides a structured paradigm for learning capabilities that are modular and interchangeable. See Appendix B for detailed related work.
@@ 3.1 Preliminary
Problem Formulation. We model the web agent’s interaction environment as a Partially Observable Markov Decision Process (POMDP), defined by the tuple $\langle\mathcal{S},\mathcal{A}_{p},\mathcal{T},\Omega,\mathcal{O}\rangle$. Here, $\mathcal{S}$ is the latent state space representing the full underlying state of the web application. $\mathcal{A}_{p}$ is the set of primitive actions the agent can execute on a webpage (e.g., click(element), type(text)), as defined in Section E.2. The function $\mathcal{T}:\mathcal{S}\times\mathcal{A}_{p}\to\Delta(\mathcal{S})$ is the stochastic state transition function. Since the agent cannot perceive the entire state $\mathcal{S}$, it receives an observation $o_{t}\in\Omega$ (e.g., the A11y tree and viewport screenshot) at each timestep $t$ through the observation function $\mathcal{O}:\mathcal{S}\to\Delta(\Omega)$.
@@ 3.2 The PolySkill framework
We identified the limitations of current skill induction methods. We tested two state-of-the-art approaches, ASI and SkillWeaver, on how well their learned skills transfer to unseen websites. Their example of skills can be seen in Table 4. As illustrated in Figure 2, our analysis revealed two key problems. First, the learning process can be unstable, producing over-specialized skills. For instance, SkillWeaver’s performance with Claude-3.7-Sonnet degrades over time because its self-proposed tasks become increasingly complex and specific. This causes the resulting skills to be too intricate and poorly suited for generalization. Second, these skills show poor generalization when applied to new websites. This is reflected in extremely low Skill Reusabilitys on unseen websites: less than 9% for ASI and less than 3% for SkillWeaver.
We build on prior work that demonstrates the robustness of representing skills as code (Wang et al., 2025; Liu et al., 2025a). To specifically address the brittleness and over-specialization, we introduce a solution inspired from software engineering: polymorphism. This allows our framework to separate a skill’s abstract goal from its concrete, site-specific implementation.
However, these skills are often tied to one specific website’s design, making them hard to reuse on other sites. Other methods, like SkillWeaver (Zheng et al., 2025), create robust skills by generating complex code from the agent’s experience. As shown in Figure 2, these skills use very specific code to find elements on a page, limiting them to only one site. This over-specialization makes the skills brittle: they work well on the original website but break easily on new sites with different layouts.
@@ 3.3 Skill Induction Process
Base of Skill Induction Pipeline Our skill induction process is built upon the robust verification pipeline established by ASI (Wang et al., 2025). In their framework, skill creation begins after a task is successfully completed using a sequence of primitive actions. An LLM-based induction module analyzes this successful trajectory to propose one or more programmatic skills that encapsulate reusable parts of the workflow (Pan et al., 2024). Before the skills are added to the library, a verification phase is done where the agent attempts to solve the same task again, this time by executing the newly generated skill. Only if this new execution is deemed as successful is the skill considered validated and added to the agent’s library for future use.
Algorithm 1 PolySkill: Polymorphic-Driven Skill Induction
3:for $t=1,\dots,N$ do
4:  Let $q_{t}$ be the current task from $\mathcal{Q}$.
9:   $\mathcal{K}_{t}\leftarrow\mathcal{K}_{t-1}\cup\mathcal{K}_{new}$ $\triangleright$ Update skill library for the next task
10:  else
11:   $\mathcal{K}_{t}\leftarrow\mathcal{K}_{t-1}$ $\triangleright$ Skill library remains unchanged on failure
12:  end if
13:end for
14:return $\mathcal{K}_{N}$
@@
Innovation via Polymorphic Skill Induction Where our method, PolySkill, improves is by integrating this process with a polymorphic skill structure. A critical preliminary step in our framework is that if the agent is operating on its first shopping website, it must first induce the high-level abstract class, AbstractShoppingSite, which provides a common ground of skills signature across shopping-related skills. Subsequently, during the skill induction phase, as the agent induces new skills on a specific site (e.g., amazon.com), it is guided to first register the corresponding function signature within the abstract class, and then define the concrete implementation within the site-specific class (e.g., AmazonWebsite, which inherits the AbstractShoppingSite). This reframes the induction prompt: rather than asking the LLM to directly induce skills, we instruct it to create a general abstract skill first and then implement the specific search method for an AmazonWebsite class. This encourages the agent to learn skills that are not just locally effective but are structurally consistent implementations of a shared, domain-wide concept.
@@
Skill Learning on Unseen Websites This polymorphic structure makes learning on new websites within a known category significantly more efficient. Imagine the agent has already formed the AbstractShoppingSite class derived from its initial interaction with amazon.com, and now visits walmart.com for the first time. It immediately recognizes Walmart as a shopping site and retrieves the abstract blueprint. This blueprint provides the agent with a clear set of exploration goals. Instead of randomly trying actions, it knows it needs to figure out how to concretely implement abstract skills like search_product and add_to_cart on this new site. Once the agent successfully searches for an item, it follows the standard induction process to create a new WalmartWebsite class, filling in the search_product method with the specific actions that worked. This guided approach accelerates learning by focusing the agent’s efforts on mastering the essential skills defined in the abstract interface. The pseudocode is shown in Algorithm 1.
@@ 3.4 Evaluation Setup
We evaluate our induction process over baseline, in two different settings: (1) Task-Defined Benchmarks In standard benchmark settings, we apply this process within controlled environments, including Mind2Web (Deng et al., 2023) and WebArena (Zhou et al., 2024a). Here, the agent is presented with a predefined curriculum of tasks. Each successful trajectory provides the validated sequence of actions needed to implement a concrete skill method. This controlled approach allows us to rigorously measure how well the polymorphic representation facilitates generalization and transfer to unseen tasks, websites, or domains. (2) Task-Free Continual Learning: To assess the ultimate goal of agent autonomy, we also apply our framework in a task-free setting, similar to settings as Voyager (Wang et al., 2023) and SkillWeaver (Zheng et al., 2025). In this scenario, the agent explores websites on its own, proposes its own goals, and induces skills from its successful attempts. Crucially, we showed that our polymorphic hierarchy enables structured exploration (Murty et al., 2025; Gandhi and Neubig, 2025). The already-learned abstract domain classes act as a schema, providing a strong prior for what skills are worth discovering (Liang et al., 2025).
@@ 3.5 Evaluation Metrics
We assess performance using five key metrics. Two are standard benchmarks adopted from prior work (Wang et al., 2024; 2025), while we introduce three new metrics to measure skills’ utility. (1) Task Success Rate (SR) is the percentage of held-out tasks the agent completes successfully. This is the primary measure of overall performance. (2) Number of Steps is the average number of actions the agent takes to complete a task, where each primitive action and each call to a skill is counted as a single step. Fewer steps indicate higher efficiency. (3) Skill Reusability measures the number skill reused in new tasks. A high rate indicates the agent learns relevant, broadly applicable skills rather than overly niche ones. (4) Task Coverage measures the tasks that used at least one skill. This metric indicates whether the skills can be adaptive in actual test case scenarios. (5) Skill Compositionality measures how frequently the system reuses existing skills as building blocks for more complex tasks. A high score indicates an efficient and scalable learning process, as the system leverages its acquired knowledge rather than learning every new skill from scratch. Formal definitions for all metrics are provided in Appendix E.1.
@@ 4 Experiments
Figure 3: Performance comparison of PolySkill with baseline methods on the Mind2Web benchmark across four large language models. The y-axis shows task success rate (%). The three evaluation settings on the x-axis, Cross-task, Cross-Website, and Cross-Domain, represent increasing levels of generalization difficulty. Our method, PolySkill, consistently outperforms the ASI baseline, while the online continual update version, PolySkill (+Update), achieves state-of-the-art performance across all models and settings. The performance gains are most significant in the challenging Cross-Domain scenario, highlighting our method’s superior generalization. Error bars represent the standard error over three runs.
@@ 4.1 Standard Benchmark Evaluation
Results. For Mind2Web, as shown in Figure 3 (Results for WebArena are presented in Table 8). Our PolySkill outperforms ASI across both static and online settings. On GPT-4.1, PolySkill improves Cross-task accuracy from 52–53% (ASI) to 55–56%, and its online variant jumps to about 63%, compared to only 59% for ASI (+Update). In the hardest Cross-Domain setting, PolySkill (+Update) reaches 63–64% versus 62% for ASI (+Update). On Qwen3-Coder, PolySkill (+Update) improves Cross-task from 41.5% (ASI) to 47.5% and Cross-Domain from 35.2% to 39.9%. These consistent gains, particularly the significant jumps in Cross-Website, underscore PolySkill ’s stronger generalization and the clear benefit of continual online updating over ASI.